OpenStudy (anonymous):

i think i kinda get this but just need a little help Find an equation for the nth term of the arithmetic sequence. 8, 6, 4, 2, ...

OpenStudy (mertsj):

What is the first term?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

i think it might be this one because its going down by 2 each time but im not toattly sure an = 8 - 2

OpenStudy (mertsj):

What is the common difference?

OpenStudy (wolf1728):

an = 8-2*n n=0 an = 8 n=1 an = 6 n=2 an = 4